Problem A.3

The formula for calculating the sum of all natural integers from 1 to n is well-known: Sn =1+2+3+...+n= n2 +n

 2

Similary, we know about the formula for calculating the sum of the first n squares:

n3 n2 n Qn =1·1+2·2+3·3+...+n·n= 3 + 2 + 6

Now, we reduce one of the two multipliers of each product by one to get the following sum: Mn =0·1+1·2+2·3+3·4+...+(n−1)·n

Find an explicit formula for calculating the sum Mn.


1
Expert's answer
2021-11-22T15:46:49-0500

Given that sum of all natural numbers 1 to n is:

"S_{n} = 1 + 2 + 3+ ..... + n = \frac{n\\^2 + n }{2} ------- (1)"


Sum of first n natural numbers square is:

"S_{n} = 1.1 + 2.2 + 3.3 + ..... + n.n \\
\frac{n\\^3}{3} + \frac{n\\^2}{2}+ \frac{n}{6}= .................. (2)"

Substract equation (1) from equation (2)


"=(1.1 + 2.2 + 3.3 + ..... + n.n) - (1 + 2 + 3+ ..... + n)"

"=(1.1 - 1+ 2.2 -2 + 3.3 -3 + ..... + n.n - n)"

"= \frac{n\\^3}{3} + \frac{n\\^2}{2}+ \frac{n}{6} - \frac{n\\^2 + n }{2}"

"=0+ 2(2 -1) + 3(3 -1) + ..... + n(n - 1)"

"=0+ 1. 2 + 2.3 + ..... + n(n - 1)"

"=\frac{ 2n\\^3 - 2n}{6}"

"=\frac{ n\\^3 - n}{3}"

Th explicit formula for calculating the sum Mn is "=\frac{ n\\^3 }{3} - \frac{n}{3}"